# Exceptional Points in a $\mathcal{PT}$-symmetrical quantum system: a Scattering matrix approach

### 05 Dec 2023 v1

@@ Acknowledgements.
J. C.-G. thanks financial support from CONAHCyT. V. D.-R is grateful with the Sistema Nacional de Investigadores, Mexico, and with E. Michel-Hernández and G. T. Domínguez-Michel for their encouragement. G. B. and V. D.-R. thank financial support for the CONAHCyT project CB2017-2018/A1-S-33920 and the UAM-Azc project CB004-22.
V Appendix
V.1 $S$ matrix building process
The first notation uses two joined subscripts. For example, in $r^{\prime}_{\eta\delta}$ we refer to the reflection that occurs between the adjacent potentials $\eta$ and $\delta$ when waves incide from the right. The scattering matrix of this case is given by
@@ V.1 $S$ matrix building process
where $k_{\delta}$ and $k_{\eta}$ are the wave numbers of the region $\delta$ and $\eta$, respectively. Also, $S_{\delta\eta}$ represents the scatterer between the $\delta$ and $\eta$ regions.
The second notation uses two subscripts separated by a hyphen. Combinating the scattering matrices $S_{\delta\eta}$ and $S_{\eta\omega}$ gives $S_{\delta-\omega}$. For example, $t_{\delta-\omega}$ refers to the total transmission generated when the waves are transmitted from region $\delta$ to the non-contiguos region $\omega$ when incident waves arrive from the left. Part of the wave incident from the left on the scatterer $S_{\delta\eta}$ is reflected in the region $\delta$ and the other part is transmitted towards the region $\eta$. The part that was transmitted travels to the scatterer $S_{\eta\omega}$, gaining a phase in this path, in which a fraction of that wave is reflected in the region $\eta$ and the other is transmitted towards the region $\omega$. The wave that was reflected travels again to the scatterer $S_{\delta\eta}$ in which a part of that wave is reflected again into the region $\eta$ and the other part is transmitted to the region $\delta$. This process of multiple reflections in the region $\eta$ is repeated infinitely. On the one hand, the sum of all the waves in the region $\omega$ is a geometric series and results in $t_{\delta-\omega}$. On the other hand, the sum of the waves in the region $\delta$, which is also a geometric series, gives us $r_{\delta-\omega}$. These two elements are the transmission and reflection coefficients, when waves are incident from the left, of the matrix
@@
This procedure is also valid if it is necessary to build the system from the right since $S=S_{1-7}=S_{7-1}$. In particular, this combination procedure is performed five times for the dimer by adding potential barriers from left to right. A part of the analytical expression of the reflection coefficient (Equation 2) of the dimer is
@@
Here $r_{12}$ is the reflection that occurs when waves are incident from the left on the boundary between region I and region II. The second term of $r$ represents the wave that was transmitted from region I to region II, it suffers a process of multiple reflections between regions II and III, and was transmitted from region II to region I. The third term represents the wave that was transmitted from region I region II, it suffers a process of multiple reflections between the boundaries of region II, was transmitted from region II to region III, was reflected multiple times between the boundaries of region I and region IV (note the term $r^{\prime}_{1-3}$ in the denominator), was transmitted from region III to region II, reflected multiple times in region II, and was finally transmitted from region II to region I. The physical interpretation continues in the same way for subsequent terms. The interpretation of the transmission coefficient follows a process equivalent to that of $r$.

### 10 Apr 2024 v2

@@ Acknowledgements.
J. C.-G. thanks financial support from CONAHCyT. V. D.-R is grateful with the Sistema Nacional de Investigadores, Mexico, and with E. Michel-Hernández and G. T. Domínguez-Michel for their encouragement. G. B. and V. D.-R. thank financial support for the CONAHCyT project CB2017-2018/A1-S-33920 and the UAM-Azc project CB004-22. The authors thank useful comments and discussions of J. A. Franco-Villafañe, A. A. Fernández-Marín and M. Martínez-Mares.
II Appendix
II.1 $S$ matrix building process
The first notation uses two joined subscripts. For example, in $r^{\prime}_{\eta\delta}$ we refer to the reflection that occurs between the adjacent potentials $\eta$ and $\delta$ when waves incide from the right. The scattering matrix for this case is given by
@@ II.1 $S$ matrix building process
where $k_{\delta}$ and $k_{\eta}$ are the wave numbers of the region $\delta$ and $\eta$, respectively, while $S_{\delta\eta}$ is the scattering matrix between the $\delta$ and $\eta$ regions which also represents the scatterer between those regions.
The second notation uses two subscripts separated by a hyphen. Combinating the scattering matrices $S_{\delta\eta}$ and $S_{\eta\omega}$ gives $S_{\delta-\omega}$. For example, $t_{\delta-\omega}$ refers to the transmission amplitude generated when the waves are transmitted from $\delta$ region to the non-contiguous $\omega$ region when incident waves arrive from the left. Part of the incident wave from the left on the scatterer $S_{\delta\eta}$ is reflected back into the $\delta$ region and the other part is transmitted towards the $\eta$ region. The part that was transmitted travels to the scatterer $S_{\eta\omega}$, gaining a phase in this path, in which a fraction of that wave is reflected in the $\eta$ region and the other is transmitted towards the $\omega$ region. The wave that was reflected travels again to the scatterer $S_{\delta\eta}$ in which a part of that wave is reflected again into the $\eta$ region and the other part is transmitted to the $\delta$ region. This process of multiple reflections in the region $\eta$ is repeated infinitely. On the one hand, the sum of all the waves in the $\omega$ region is a geometric series which results in $t_{\delta-\omega}$. On the other hand, the sum of the waves in the $\delta$ region, which is also a geometric series, gives $r_{\delta-\omega}$. These two elements are the transmission and reflection amplitudes when waves are incident from the left of the matrix
@@
This procedure is also valid if it is necessary to build the system from the right since $S=S_{1-7}=S_{7-1}$. In particular, this combination procedure is performed five times for the dimer by adding potential barriers from left to right. A part of the analytical expression of the reflection amplitude (Equation 2) of the dimer is
@@
Here $r_{12}$ is the first reflection that occurs when waves are incident from the left on the boundary between regions I and II. The second term of $r$ represents the wave that was transmitted from region I to region II, it suffers a process of multiple reflections between regions II and III, and is transmitted from region II to region I. The third term represents the wave that is transmitted from region I to region II, it suffers a process of multiple reflections between the boundaries of region II, is transmitted from region II to region III, is reflected multiple times between the boundaries of region I and region IV (note the term $r^{\prime}_{1-3}$ in the denominator), is transmitted from region III to region II, reflected multiple times in region II, and is finally transmitted from region II to region I. The physical interpretation continues in the same way for subsequent terms. The interpretation of the transmission amplitude follows an equivalent process to that of $r$.